xy' - (x+2)y = 0, x0 = 0


1
Expert's answer
2021-05-07T09:03:38-0400

"xy' - (x+2)y = 0"

separate variables : "\frac{dy}{y} = \frac{x+2}{x} dx \Rightarrow d(ln(y)) = (1 + \frac{2}{x})dx"

integrate : "ln(y) = x + 2\cdot ln(x) + const" , "y = x^2 \cdot e^{x} \cdot C" , where C = const


ANSWER : "y = x^2 \cdot e^{x} \cdot C" , where C = const